# Do MLLMs Really Understand the Charts?

### v1

@@ Conclusion
    We cloned the official public repository associated with each chart-specific model.
@@
    We fed the images and questions from our CRBench test set into these scripts.
This methodology ensures that the results reported in the main body of the manuscript represent each model performing under its own optimal, author-intended conditions. This prevents any potential performance degradation that might arise from imposing a foreign prompt structure and guarantees that we are comparing ChartReasoner against the most capable version of each specialized model.
Evaluation of ChartReasoner on Public Benchmarks
To validate the generalization of ChartReasoner’s enhanced reasoning capabilities, we evaluated it against several standard public benchmarks. The following subsections detail the exact procedure used for each.
CharXiv
For the CharXiv benchmark, we utilized the official code and evaluation scripts available in their repository. Our evaluation process involved adapting their provided inference pipeline. We integrated our locally-deployed ChartReasoner (served via vLLM) as the model backend, replacing the default model call. All other components, including data loading, pre-processing, and the final scoring scripts for both the reasoning (R) and descriptive (D) tasks, were kept identical to the original setup.
ChartBench
Our evaluation on ChartBench followed a similar protocol. We used the complete evaluation pipeline provided in the official ChartBench repository. Predictions for each test instance were generated by passing the questions to our ChartReasoner vLLM endpoint. The resulting outputs were then fed directly into the official scoring script, which computes the overall accuracy based on their defined metrics and question categories.
ChartQAPro
The official repository for ChartQAPro provides a standalone evaluation script but does not include a full inference pipeline. Therefore, to benchmark ChartReasoner, we implemented a two-step process that carefully reconstructs the intended evaluation conditions: 1. Prediction Generation: We developed a custom script to generate predictions for the entire ChartQAPro test set. For each question, this script formatted a query for ChartReasoner that precisely replicated the direct prompt template described in the official ChartQAPro paper. 2. Scoring: The complete set of generated predictions was saved to a file. This file was then used as input for the official evaluation script provided by the ChartQAPro authors. The script automatically compares our model’s predictions against the ground-truth annotations to compute the final accuracy score.
Figure 6: An overview of sample question-answer pairs for various synthetic chart types within the CRBench dataset.
Figure 7: An overview of sample question-answer pairs for various real chart types within the CRBench dataset.
@@
Figure 11: Direct Prompting in CRBench Evaluation
Figure 12: Prompting for Optional Output of Reasoning Steps in CRBench Evaluation
Figure 13: Prompting for Forced Output of Reasoning Steps in CRBench Evaluation
Figure 14: Example 1 from CRBench comparing the CoT outputs of Qwen2.5-VL-7B and ChartReasoner-7B.
Figure 15: Example 2 from CRBench comparing the CoT outputs of Qwen2.5-VL-7B and ChartReasoner-7B.

### v2

@@ C.2 Evaluation on ChartVRBench
    We cloned the official public repository for each model.
@@
    We fed the images and questions from our ChartVRBench test set directly into these scripts.
This methodology ensures that we are comparing our model against the most capable version of each specialized baseline.
C.3 Evaluation of ChartVR on Public Benchmarks
To validate the generalization of ChartVR’s enhanced reasoning capabilities, we evaluated it against several standard public benchmarks, following the official protocol for each.
CharXiv.
We utilized the official code and evaluation scripts from the CharXiv repository. We integrated our locally-deployed ChartVR as the model backend into their inference pipeline, keeping all other components (data loading, pre-processing, and scoring scripts) identical to the original setup.
ChartBench.
Our evaluation followed a similar protocol using the complete pipeline from the official ChartBench repository. We generated predictions with our ChartVR model and fed the outputs directly into the official scoring script.
ChartQAPro.
As the official repository provides a standalone evaluation script but not a full inference pipeline, we implemented a two-step process. First, we developed a script to generate predictions for the test set using a prompt that precisely replicated the template described in the ChartQAPro paper. Second, the resulting file of predictions was used as input for the official evaluation script to compute the final accuracy score.
Appendix D Interpretation of Human Performance
It is important to note that human accuracy on this task is not 100%. This is primarily due to inherent tendencies in human visual estimation, for instance, individuals often gravitate towards estimating with round or integer values that appear close to the correct answer, rather than performing precise interpolation. Our analysis indicates that a 2% relative error tolerance is a reasonable threshold to account for these natural human inaccuracies.
Furthermore, performance varies significantly across chart types. For area charts, accuracy sees a substantial decline, because many questions require calculating the difference between the upper and lower boundaries of a shaded region, a task made considerably more difficult by the common absence of horizontal gridlines as visual aids. Similarly, for complex combo charts, lower performance can be attributed to cognitive factors, such as misinterpretation of the prompt or misunderstanding the intricate relationships between different chart components.
Figure 4: An overview of sample question-answer pairs for various synthetic chart types within the ChartVRBench dataset.
Figure 5: An overview of sample question-answer pairs for various complex synthetic chart types within the ChartVRBench dataset.
Figure 6: An overview of sample question-answer pairs for various real chart types within the ChartVRBench dataset.
@@ Appendix D Interpretation of Human Performance
Figure 10: CoT Prompting Reasoning Steps in ChartVRBench Evaluation
Figure 11: Example 1 from ChartVRBench comparing the CoT outputs of Qwen2.5-VL-7B and ChartVR-7B.
Figure 12: Example 2 from ChartVRBench comparing the CoT outputs of Qwen2.5-VL-7B and ChartVR-7B.